# Conserved charges and asymptotic symmetries of BTZ-like black holes in Einstein-bumblebee gravity

### v1

@@ I Introduction
Symmetry is fundamental in modern physics. According to the well known Noether’s theorem, symmetries relate to conservation laws Noether:1918zz . In the realm of black hole physics, the analysis of the asymptotic symmetry of spacetime geometry under appropriate boundary conditions brings new comprehension and great importance. The investigation of asymptotic symmetries in gravitational theories started from the works of Bondi-Metzner-Sachs (BMS) in 1962 Bondi:1962px ; Sachs:1962wk , which identified the BMS group of supertranslations and Lorentz transformations as asymptotic symmetry group (ASG). Another landmark discovery by Brown and Henneaux Brown:1986nw established a definitive link between the asymptotic symmetries of three-dimensional asymptotic anti-de-Sitter ($\mathrm{AdS_{3}}$) spacetime and the symmetries of two-dimensional conformal field theory (CFT), serving as an early precursor of the AdS/CFT correspondence Maldacena:1998bw . By applying the AdS/CFT to the Banados-Teitelboim-Zanelli (BTZ) black holes Banados:1992wn ; Banados:1992gq , the microscopic origin of the Bekenstein-Hawking entropy was derived Strominger:1997eq . A parallel generalization of the AdS/CFT correspondence in near horizon region is the Kerr/CFT correspondence Guica:2008mu , which links near horizon geometry of a four-dimensional extremal Kerr black hole with a two-dimensional chiral CFT.
@@
In the theories of gravity, conserved charges such as black hole mass, angular momentum, electric charge and black hole entropy associated with exact and asymptotic symmetries have been extensively studied by various methods (see Compere:2018aar ; Adami:2017phg ; Frodden:2019ylc for recent reviews of the methods for conserved charges). Based on the Noether procedure, the covariant phase space method (CPSM) was proposed by Wald and Iyer (also called Wald’s formalism) Lee:1990nz ; Wald:1993nt ; Iyer:1994ys . In the framework of CPSM, a systematic approach for calculating conserved charges in generally covariant theories of gravity was developed by K. Hajian et al, which is called solution phase space method (SPSM) Hajian:2015xlp . The SPSM is a restriction of CPSM to solution space manifold, focusing on parametric variation and exact symmetries. In SPSM, the charge variations are computed by an integration over an arbitrary smooth, closed, and compact codimension-2 surface surrounding the black hole singularity. For reviews and applications of SPSM see Refs.Hajian:2016kxx ; Ghodrati:2016vvf ; Ding:2020bwa ; Guo:2024zaa . The study of conserved charges and asymptotic symmetries is crucial for understanding the thermodynamic properties and holographic dualities in gravitational theories. Generally, a physical system is governed by its action and the boundary conditions; the former determines the equations of motion (and therefore the dynamics of the system). This also implies that different actions and boundary conditions contain different conserved charges and ASG.
@@
As a special family of modified gravitational theories, the Lorentz-violating (LV) theories of gravity Mattingly:2005re , including Hořava-Lifshitz theory Horava:2009uw , ghost condensation Arkani-Hamed:2003pdi , warped brane world Frey:2003jq ; Cline:2003xy , Einstein-aether theory Jacobson:2000xp ; Eling:2004dk ; Jacobson:2007veq and Einstein bumblebee gravity Kostelecky:1988zi ; Kostelecky:1989jw ; Kostelecky:2003fs ; Kostelecky:2010ze , have attracted increasing interests. In the action of the bumblebee gravity, the spontaneous Lorentz violation arises from a potential $V(B_{\mu}B^{\mu})$ acting on a bumblebee vector field $B_{\mu}$ Kostelecky:2003fs ; Kostelecky:2010ze . In 2017, Casana et al obtained an exact Schwarzschild-like black hole solution in bumblebee gravity Casana:2017jkc . Subsequently, a number of exact solutions are obtained, including Schwarzschild-anti-de-Sitter (AdS)-like black holes Maluf:2020kgf , slowly rotating Kerr-like black holes Ding:2019mal , rotating BTZ-like black holes Ding:2023niy , and other solutions Santos:2014nxm ; Jha:2020pvk ; Filho:2022yrk ; Xu:2022frb ; Liu:2024axg within bumblebee gravity, all of which exhibit effects induced by LV . Therefore, whether the LV effects in the bumblebee gravity influence the conserved charges and ASG remains an open question worthy of further study.
@@
Although the Lorentz symmetry is broken in the theories of bumblebee gravity, it remains generally invariant under diffeomorphism transformation Bluhm:2007bd . Therefore, the SPSM is still accommodation. In this paper, we employ the SPSM to investigate the conserved charges of a rotating BTZ-like black hole Ding:2023niy in Einstein-bumblebee gravity. We compute the black hole mass, angular momentum and entropy, showing that these quantities are affected by LV parameter. As a consistent check, we also derive the first law of black hole thermodynamics. Mostly, the ASG analysis is carried out in pure gravity Compere:2008cv ; Compere:2009zj ; Blagojevic:2009ek ; Detournay:2023zni ; Godet:2020xpk . In this work, we investigate the ASG in the gravity coupled with the bumblebee field non-minimally. Given that the black hole solution describes an asymptotic AdS spacetime, we study the AdS/CFT correspondence and obtain the asymptotic charge algebra, which consists of two copies of the Virasoro algebra with non-trivial central charges. Using the Cardy formula Cardy:1986ie , we compute the microscopic entropy of dual CFT and find that it is in precise agreement with the Bekenstein-Hawking entropy. Recently, a new class of boundary conditions was devised by Detournay-Smoes-Wutte to study asymptotic symmetries in near horizon geometry Detournay:2023zni , which is a higher dimensional uplift of Godet-Marteau boundary conditions in two-dimensional gravity Godet:2020xpk . An application of these new boundary conditions is given in Ref.Jiang:2024tmh . In near horizon region, we apply the Detournay-Smoes-Wutte boundary conditions to study the asymptotic symmetries for near horizon geometry of extremal BTZ-like black hole. As a result, we obtain a Virasoro-Kac-Moody $\mathrm{U(1)}$ algebra, which is the symmetry algebra of a warped conformal field theory (WCFT).
@@
This paper is organized as follows. In Sec.2, we provide a brief review of the SPSM. In Sec.3, we derive the Noether-Wald charge density and the surface charge density for Einstein-bumblebee gravity. In Sec.4, we calculate the conserved charges and check the first law of thermodynamics for rotating BTZ-like black hole. In Sec.5, we study the asymptotic symmetries in asymptotic $\mathrm{AdS_{3}}$ background. Two copies of Virasoro charge algebra will be obtained. Subsequently, we derive the microscopic entropy of dual CFT. In the near horizon geometry of extremal BTZ-like black hole, we apply the new boundary conditions to obtain the symmetry algebra of WCFT. Finally, Sec.6 concludes with a summary of our main findings and remarks.
II Solution phase space method
@@
The SPSM Hajian:2015xlp is based on the CPSM Lee:1990nz ; Wald:1993nt ; Iyer:1994ys , so firstly we give a brief review of CPSM to achieve the conserved charges for generic gauge theories.
@@
where $\delta\Phi$ is a generic field variation and provides the basis for tangent bundle of the phase space $\mathcal{M}$. The set $\mathbf{E}_{\Phi}=0$ gives the equations of motion (EOM). $\mathbf{\Theta}$ is an $(n-1;1)$-form called Lee-Wald symplectic potential, where the $(n-1)$-form is over the spacetime and the $1$-form is on the tangent bundle of $\mathcal{M}$. The Lee-Wald symplectic form Lee:1990nz is defined as
@@
representing the symplectic current form, which is an $(n-1;2)$-form. $\Sigma$ is a smooth codimension-$1$ surface.
@@
Here $\bm{k}_{\epsilon}$ is an $(n-2;1)$-form, called the surface charge density. The $(n-2;2)$-form $\mathbf{Q}_{\epsilon}$ is Noether-Wald charge density, related to the Noether current $\mathbf{J}_{\epsilon}$ by $\mathbf{J}_{\epsilon}\approx\mathrm{d}\mathbf{Q}_{\epsilon}$, and
@@
In Eq.(II), $\delta^{[\Phi]}$ denotes that $\delta$ acts only on $\Phi$ but not on the $\epsilon$ inside $\mathbf{\Theta}$.
If $\delta H_{\epsilon}(\Phi)$ is well-defined and integrable, we could determine the conserved charge $H_{\epsilon}$. The integrability condition is basically $\left(\delta_{1}\delta_{2}-\delta_{2}\delta_{1}\right)H_{\epsilon}(\Phi)=0$, which is equivalent to Hajian:2015xlp
@@
The corresponding symplectic $2$-form $\Omega$ is denoted as $\hat{\Omega}$. The solution phase space is denoted by (^$(\hat{\mathcal{M}},\hat{\Omega})$, which is a submanifold of $(\mathcal{M},\Omega)$ when $\{\Phi\}$ are restricted to solution space $\{\hat{\Phi}\}$. The charge variations associated with the generators $\epsilon$ are taken as
$\hat{\delta}H_{\epsilon}(\Phi)=\oint_{\partial\Sigma}\bm{k}_{\epsilon}(\hat{\delta}\Phi,\hat{\Phi}).$ (13)
@@
Algebra of surface charges: When the charge variation $\delta H_{\epsilon}(\Phi)$ is well-defined and integrable, the algebra of the asymptotic symmetries is given by the Poisson bracket algebra of the surface charges Barnich:2007bf ; Compere:2007az
@@
where $[\epsilon,\tilde{\epsilon}]$ denotes the Lie bracket of the generator $\epsilon$, and the last term $i\oint_{\partial\Sigma}\bm{k}_{\epsilon}(\delta_{\tilde{\epsilon}}\bar{\Phi},\bar{\Phi})$ is the central extention term in the algebra.
III Einstein-bumblebee theory
In the bumblebee theory of gravity, the bumblebee vector field $B_{\mu}$ is introduced to couples with gravity non-minimally. This field acquires a nonzero vacuum expectation value, leading to the spontaneous breaking of Lorentz symmetry in the gravitational sector. For Einstein-bumblebee gravity, the dynamical fields $\Phi$ consist of the metric $g_{\mu\nu}$ and bumblebee field $B_{\mu}$. The Lagrangian is given by Ding:2023niy
@@ III Einstein-bumblebee theory
The potential $V(B_{\mu}B^{\mu}\pm b^{2})$ triggers Lorentz violation, leading to a nonvanishing vacuum expectation value for $B_{\mu}$, such that $B^{\mu}\right>=b^{\mu}$, where the vector $b_{\mu}$ satisfies $b^{\mu}b_{\mu}=\mp b^{2}=const$.
@@
$\delta\mathbf{L}(\Phi)=\left(\mathrm{E}^{\mu\nu}_{g}\delta g_{\mu\nu}+\mathrm{E}^{\mu}_{B}\delta B_{\mu}\right)\bm{\epsilon}+\mathrm{d}\mathbf{\Theta}(\delta\Phi,\Phi)$ (20)
@@
$k^{\mu\nu}_{gB}(\delta\Phi,\Phi)=$ $\frac{\rho}{2\kappa}\cdot 2\bigg{[}\frac{1}{2}h\nabla^{[\mu}(B^{\nu]}B_{\sigma})\xi^{\sigma}+h\nabla_{\sigma}(B^{\sigma}B^{[\mu})\xi^{\nu]}-\frac{1}{2}h\nabla_{\sigma}(B^{\sigma}B^{[\mu}\xi^{\nu]})-h^{\alpha[\mu}\nabla_{\alpha}(B^{\nu]}B_{\sigma})\xi^{\sigma}$
$+g^{\alpha[\mu}\delta[\nabla_{\alpha}(B^{\nu]}B_{\sigma})]\xi^{\sigma}+2\delta[\nabla_{\sigma}(B^{\sigma}B^{[\mu})]\xi^{\nu]}-\delta[\nabla_{\sigma}(B^{\sigma}B^{[\mu}\xi^{\nu]})]$
$-[\nabla_{\alpha}(\Omega^{[\mu|\alpha\sigma\rho|}h_{\sigma\rho})+\Sigma^{[\mu|\alpha\sigma|}h_{\alpha\sigma}]\cdot\xi^{\nu]}\bigg{]},$
@@
IV Conserved charges for BTZ-like bumblebee black hole
When the potential $V=\frac{\lambda}{2}X$ in the Lagrangian, where $\lambda$ is a Lagrange-multiplier, and there is no bumblebee field frosted at its vacuum expectation value, i.e., $B_{\mu}=b_{\mu}$, the $3$-dimensional rotating BTZ-like bumblebee black hole solution is given by Ding:2023niy

### v2

@@ 1 Introduction
Symmetry is fundamental in modern physics. According to the well-known Noether’s theorem, symmetries are related to conservation laws [1]. In the realm of black hole physics, the analysis of the asymptotic symmetry of spacetime geometry under appropriate boundary conditions brings new comprehension and great importance. The investigation of asymptotic symmetries in gravitational theories started from the works of Bondi-Metzner-Sachs (BMS) in 1962 [2, 3], which identified the BMS group of supertranslations and Lorentz transformations as asymptotic symmetry group (ASG). Another landmark discovery by Brown and Henneaux [4] established a definitive link between the asymptotic symmetries of three-dimensional asymptotic anti-de-Sitter ($\mathrm{AdS_{3}}$) spacetime and the symmetries of two-dimensional conformal field theory (CFT), serving as an early precursor of the AdS/CFT correspondence [5]. By applying the AdS/CFT to the Banados-Teitelboim-Zanelli (BTZ) black holes [6, 7], the microscopic origin of the Bekenstein-Hawking entropy was derived [8]. A parallel generalization of the AdS/CFT correspondence in near horizon region is the Kerr/CFT correspondence [9], which links near horizon geometry of a four-dimensional extremal Kerr black hole with a two-dimensional chiral CFT. Another holographic realization beyond AdS/CFT is the $\mathrm{WAdS_{3}/WCFT}$ correspondence, relating warped $\mathrm{AdS_{3}}$ ($\mathrm{WAdS_{3}}$) to warped CFT (WCFT) [10].
@@
In the theories of gravity, conserved charges such as black hole mass, angular momentum, electric charge and black hole entropy associated with exact and asymptotic symmetries have been extensively studied by various methods (see [11, 12, 13] for recent reviews of the methods for conserved charges). Based on the Noether procedure, the covariant phase space method (CPSM) was proposed by Wald and Iyer (also called Wald’s formalism) [14, 15, 16]. In the framework of CPSM, a systematic approach for calculating conserved charges in generally covariant theories of gravity was developed by K. Hajian et al, which is called solution phase space method (SPSM) [17]. The SPSM is a restriction of CPSM to solution space manifold, focusing on parametric variation and exact symmetries. In SPSM, the charge variations are computed by an integration over an arbitrary smooth, closed, and compact codimension-2 surface surrounding the black hole singularity. For reviews and applications of SPSM see Refs.[18, 19, 20, 21]. The study of conserved charges and asymptotic symmetries is crucial for understanding the thermodynamic properties and holographic dualities in gravitational theories. Generally, a physical system is governed by its action and the boundary conditions; the former determines the equations of motion (and therefore the dynamics of the system). This also implies that different actions and boundary conditions contain different conserved charges and ASG.
@@
As a special family of modified gravitational theories, the Lorentz-violating (LV) theories of gravity [22], including Hořava-Lifshitz theory [23], ghost condensation [24], warped brane world [25, 26], Einstein-aether theory [27, 28, 29] and Einstein bumblebee gravity [30, 31, 32, 33], have attracted increasing interests. In the action of the bumblebee gravity, the spontaneous Lorentz violation arises from a potential $V(B_{\mu}B^{\mu})$ acting on a bumblebee vector field $B_{\mu}$ [32, 33]. In 2017, Casana et al obtained an exact Schwarzschild-like black hole solution in bumblebee gravity [34]. Subsequently, a number of exact solutions are obtained, including Schwarzschild-anti-de-Sitter (AdS)-like black holes [35], slowly rotating Kerr-like black holes [36], rotating BTZ-like black holes [37], and other solutions [38, 39, 40, 41, 42] within bumblebee gravity. Comprehensive analyses of the properties of these black holes have conducted, including gravitational lensing [43], quasinormal modes [44], Hawking radiation [45], black hole shadows [36, 46, 47, 48, 49], accretion disk [50], superradiant instability [51], parameters constraint [52, 53]. All of these properties exhibit effects induced by LV. Therefore, whether the LV effects in the bumblebee gravity influence the conserved charges and ASG remains an open question worthy of further study.
@@
Although the Lorentz symmetry is broken in the theories of bumblebee gravity, it remains generally invariant under diffeomorphism transformation [54]. Therefore, the SPSM is still accommodation. In this paper, we employ the SPSM to investigate the conserved charges of a rotating BTZ-like black hole in Einstein-bumblebee gravity [37]. We compute the black hole mass, angular momentum and entropy, showing that these quantities are affected by LV parameter. As a consistent check, we also derive the first law of black hole thermodynamics. Mostly, the ASG analysis is carried out in pure gravity [55, 56, 57, 58, 59]. In this work, we investigate the ASG in the gravity coupled with the bumblebee field non-minimally. Given that the black hole solution describes an asymptotic AdS spacetime, we study the AdS/CFT correspondence and obtain the asymptotic charge algebra, which consists of two copies of the Virasoro algebra with non-trivial central charges. Using the Cardy formula [60], we compute the microscopic entropy of dual CFT and find that it is in precise agreement with the Bekenstein-Hawking entropy. It is of interest to expand the scope of holography beyond asymptotically $\mathrm{AdS}$ spacetime. The most notable non-$\mathrm{AdS}$ backgrounds include $\mathrm{WAdS_{3}}$ spacetimes and near horizon geometry of extremal Kerr black holes (NHEK), these geometries have $SL(2,\mathbb{R})\times U(1)$ isometries [61, 62]. The boundary conditions can be chosen so that the symmetries are enhanced to a semi-direct sum of Virasoro algebra and Kac-Moody $\mathrm{U(1)}$ algebra, which is the symmetry algebra of WCFT [10], suggesting that WCFT could be a holographic counterpart to $\mathrm{WAdS_{3}}$ or NHEK [55, 56, 57, 58, 63, 64, 65]. Besides being characterized by Virasoro-Kac-Moody $\mathrm{U(1)}$ algebra, WCFT is a two-dimensional non-relativistic quantum field theory (QFT) that exhibits Lorentz symmetry breaking. Therefore, it is important to explore a potential dual holographic model which violates Lorentz symmetry on both the gravity and the field theory sides, as well as to investigate the influence of LV effects on the structure of ASG and the central charges. Recently, a new class of boundary conditions was devised by Detournay-Smoes-Wutte to study asymptotic symmetries in near horizon geometry [58], which is a higher dimensional uplift of Godet-Marteau boundary conditions in two-dimensional gravity [59]. An application of these new boundary conditions is given in Ref.[66]. In near horizon region, we apply the Detournay-Smoes-Wutte boundary conditions to study the asymptotic symmetries for near horizon geometry of extremal BTZ-like black hole. As a result, we obtain a Virasoro-Kac-Moody $\mathrm{U(1)}$ algebra, this provides a first step towards building a holographic dual for near horizon geometry of bumblebee black holes.
@@
This paper is organized as follows. In Sec.2, we provide a brief review of the SPSM. In Sec.3, we derive the Noether-Wald charge density and the surface charge density for Einstein-bumblebee gravity. In Sec.4, we calculate the conserved charges and check the first law of thermodynamics for rotating BTZ-like black hole. In Sec.5, we study the asymptotic symmetries in asymptotic $\mathrm{AdS_{3}}$ background. Two copies of Virasoro charge algebra will be obtained. Subsequently, we derive the microscopic entropy of dual CFT. In the near horizon geometry of extremal BTZ-like black hole, we apply the new boundary conditions to obtain the symmetry algebra of WCFT. Finally, Sec.6 concludes with a summary of our main findings and outlooks.
2 Solution phase space method
@@
The SPSM [17] is based on the CPSM [14, 15, 16], so firstly we give a brief review of CPSM to achieve the conserved charges for generic gauge theories.
@@
where $\delta\Phi$ is a generic field variation and provides the basis for tangent bundle of the phase space $\mathcal{M}$. The set $\mathbf{E}_{\Phi}=0$ gives the equations of motion (EOM). $\mathbf{\Theta}$ is an (n1;1)(n-1;1)-form called Lee-Wald symplectic potential, where the (n1)(n-1)-form is over the spacetime and the $1$-form is on the tangent bundle of $\mathcal{M}$. The Lee-Wald symplectic form [14] is defined as
@@
representing the symplectic current form, which is an (n1;2)(n-1;2)-form. $\Sigma$ is a smooth codimension-$1$ surface.
@@
Here $\bm{k}_{\epsilon}$ is an (n2;1)(n-2;1)-form, called the surface charge density. The (n2;2)(n-2;2)-form $\mathbf{Q}_{\epsilon}$ is Noether-Wald charge density, related to the Noether current $\mathbf{J}_{\epsilon}$ by $\mathbf{J}_{\epsilon}\approx\mathrm{d}\mathbf{Q}_{\epsilon}$, and
@@
In Eq.(2), $\delta^{[\Phi]}$ denotes that $\delta$ acts only on $\Phi$ but not on the $\epsilon$ inside $\bm{\Theta}$.
If $\delta H_{\epsilon}(\Phi)$ is well-defined and integrable, we could determine the conserved charge $H_{\epsilon}$. The integrability condition is basically $\left(\delta_{1}\delta_{2}-\delta_{2}\delta_{1}\right)H_{\epsilon}(\Phi)=0$, which is equivalent to [17]
@@
[…]
@@
Algebra of surface charges: When the charge variation $\delta H_{\epsilon}(\Phi)$ is well-defined and integrable, the algebra of the asymptotic symmetries is given by the Poisson bracket algebra of the surface charges [67, 68]
@@
where $\epsilon,\tilde{\epsilon}$] denotes the Lie bracket of the generator $\epsilon$, $i$ is the imaginary unit, and the last term $i\oint_{\partial\Sigma}\bm{k}_{\epsilon}(\delta_{\tilde{\epsilon}}\bar{\Phi},\bar{\Phi})$ is the central extention term in the algebra.
3 Einstein-bumblebee theory
In the bumblebee theory of gravity, bumblebee vector field $B_{\mu}$ is introduced to couples with gravity non-minimally. This field acquires a nonzero vacuum expectation value, leading to the spontaneous breaking of Lorentz symmetry in the gravitational sector. For Einstein-bumblebee gravity, the dynamical fields $\Phi$ consist of the metric $g_{\mu\nu}$ and bumblebee field $B_{\mu}$. The Lagrangian is given by [37]
@@ 3 Einstein-bumblebee theory
The potential $V(B_{\mu}B^{\mu}\pm b^{2})$ triggers Lorentz violation, leading to a nonvanishing vacuum expectation value for $B_{\mu}$, such that $\left<B^{\mu}\right>=b^{\mu}$, where the vector $b_{\mu}$ satisfies $b^{\mu}b_{\mu}=\mp b^{2}=const$.
@@
$\delta\mathbf{L}(\Phi)=\left(\mathrm{E}^{\mu\nu}_{g}\delta g_{\mu\nu}+\mathrm{E}^{\mu}_{B}\delta B_{\mu}\right)\bm{\epsilon}+\mathrm{d}\mathbf{\Theta}(\delta\Phi,\Phi),$ (20)
@@
$k^{\mu\nu}_{gB}(\delta\Phi,\Phi)=$ $\frac{\rho}{2\kappa}\cdot 2\bigg{[}\frac{1}{2}h\nabla^{[\mu}(B^{\nu]}B_{\sigma})\xi^{\sigma}+h\nabla_{\sigma}(B^{\sigma}B^{[\mu})\xi^{\nu]}-\frac{1}{2}h\nabla_{\sigma}(B^{\sigma}B^{[\mu}\xi^{\nu]})$
$-h^{\alpha[\mu}\nabla_{\alpha}(B^{\nu]}B_{\sigma})\xi^{\sigma}+g^{\alpha[\mu}\delta[\nabla_{\alpha}(B^{\nu]}B_{\sigma})]\xi^{\sigma}+2\delta[\nabla_{\sigma}(B^{\sigma}B^{[\mu})]\xi^{\nu]}$
$-\delta[\nabla_{\sigma}(B^{\sigma}B^{[\mu}\xi^{\nu]})]-[\nabla_{\alpha}(\Omega^{[\mu|\alpha\sigma\rho|}h_{\sigma\rho})+\Sigma^{[\mu|\alpha\sigma|}h_{\alpha\sigma}]\cdot\xi^{\nu]}\bigg{]},$
@@
4 Conserved charges for BTZ-like bumblebee black hole
When the potential $V=\frac{\lambda}{2}X$ in the Lagrangian, where $\lambda$ is a Lagrange-multiplier, and there is no bumblebee field frosted at its vacuum expectation value, i.e., $B_{\mu}=b_{\mu}$, the $3$-dimensional rotating BTZ-like bumblebee black hole solution is given by [37]